# Towards a Formal Theory of the Need for Competence via Computational Intrinsic Motivation

### v1

@@ Abstract
Computational models offer powerful tools for formalising psychological theories, making them both testable and applicable in digital contexts. However, they remain little used in the study of motivation within psychology. We focus on the “need for competence”, postulated as a key basic human need within Self-Determination Theory (SDT)—arguably the most influential psychological framework for studying intrinsic motivation (IM). The need for competence is treated as a single construct across SDT texts. Yet, recent research has identified multiple, ambiguously defined facets of competence in SDT. We propose that these inconsistencies may be alleviated by drawing on computational models from the field of artificial intelligence, specifically from the domain of reinforcement learning (RL). By aligning the aforementioned facets of competence—effectance, skill use, task performance, and capacity growth—with existing RL formalisms, we provide a foundation for advancing competence-related theory in SDT and motivational psychology more broadly. The formalisms reveal underlying preconditions that SDT fails to make explicit, demonstrating how computational models can improve our understanding of IM. Additionally, our work can support a cycle of theory development by inspiring new computational models formalising aspects of the theory, which can then be tested empirically to refine the theory. While our research lays a promising foundation, empirical studies of these models in both humans and machines are needed, inviting collaboration across disciplines.
Keywords: need for competence; self-determination theory; intrinsic motivation; motivational psychology; theory development; intrinsically motivated reinforcement learning; computational modelling
@@ Introduction
Self-Determination Theory (SDT) \@BBOPcitep\@BAP\@BBN(Deci and Ryan,, 1985)\@BBCP is one of the most influential psychological theories of motivation, including intrinsic motivation (IM). It has also inspired artificial intelligence (AI) research on computational IM (e.g., \@BBOPcitep\@BAP\@BBNOudeyer and Kaplan,, 2007\@BBCP). Computational IM equips artificial agents with abilities for open-ended development, task-agnostic learning, and dealing with the potential sparsity of rewards \@BBOPcitep\@BAP\@BBN(Colas et al.,, 2022, p. 1161)\@BBCP. Notably, SDT, like many psychological theories, is expressed in “soft” verbal propositions. In light of the current theory crisis \@BBOPcitep\@BAP\@BBN(Oberauer and Lewandowsky,, 2019)\@BBCP, verbal theory has drawn increasing critique: theories that are not formally specified may fail to produce hypotheses that can be rigorously tested. While many psychology researchers have advocated for computational modelling as a paradigm for theory development (e.g., \@BBOPcitep\@BAP\@BBNMarsella et al.,, 2010; Oberauer and Lewandowsky,, 2019; Robinaugh et al.,, 2021\@BBCP), formal specification in the form of precisely defined computational models remains largely absent in SDT. Judging by the most recent integrative volume of the theory, the Oxford Handbook of Self-Determination Theory \@BBOPcitep\@BAP\@BBN(Ryan,, 2023)\@BBCP, methodological advances chiefly focus on new psychometric approaches and neuroscientific evidence.
Here, we make the case that computational models of IM can benefit the development of SDT. More specifically, \@BBOPcitet\@BAP\@BBNDeterding et al., (2024)\@BBCP have shown that the need for competence, one of the basic psychological needs underpinning IM in SDT, is poorly specified, with multiple definitions that lack clear operationalisation. They derive four distinct facets of competence by means of a detailed analysis of SDT texts spanning many decades (\@BBOPciteyear\@BAP\@BBN1985\@BBCP\@BBOPciteyear\@BAP\@BBN2023\@BBCP), reproduced in Table 1 \@BBOPcitep\@BAP\@BBN(Deterding et al.,, 2024, pp. 7–10)\@BBCP. We focus on matching computational formalisms to each of these facets.
@@
Task performance (C3) Observing that one performs (a) well at an intended task, (b) to an extent that requires a certain skill or skill level.
@@
One reason computational models of IM offer a powerful paradigm for developing SDT is that translating a concept like competence into an algorithmic description induces a different way of thinking about theory: the process forces us to specify abstract and informal verbal theories to the point that they can be turned into a system that can actually produce observable behaviour. Observing how a theory behaves can “reveal implicit assumptions and hidden complexities” in verbal theory, thus inspiring theory construction \@BBOPcitep\@BAP\@BBN(Marsella et al.,, 2010, p. 4)\@BBCP. Murayama’s (\@BBOPciteyear\@BAP\@BBN2022\@BBCP) recent adoption of reinforcement learning (RL) as an underpinning framework for a theory of curiosity is a good case in point, as is Ady et al.’s work on specific machine curiosity (\@BBOPciteyear\@BAP\@BBN2022\@BBCP), to name a few.
To advance formal development of SDT and competence-related theory in motivational psychology more broadly, a key challenge is identifying formalisms that sufficiently approximate the theory. In this paper, we align each distinct competence facet with examples of formalisms from the computational literature. We then discuss how each example formalises aspects of the theory. These formalisms not only lay a foundation for specifying the need for competence but also unveil a wealth of existing implementations, simulators, and evaluation metrics from the AI literature for use in psychological motivation research. In future work, these will allow us to test whether formalisations fitting the facets, either individually or in conjunction, produce empirical results that align with propositions made by the theory.
@@
We begin by justifying why RL is a suitable domain from which to draw computational models of competence. Additionally, to support understanding of the models we discuss, we introduce some key concepts from intrinsically motivated RL—most notably, goals and skills and their relationship to the reward function—and recommend further reading.
The idea that IM might be driven by biological reward systems (\@BBOPcitep\@BAP\@BBNBarto et al.,, 2004, p. 113\@BBCP, in reference to \@BBOPcitep\@BAP\@BBNKakade and Dayan,, 2002; Dayan and Balleine,, 2002\@BBCP) has driven the choice of methods centrally used in computational IM: RL \@BBOPcitep\@BAP\@BBN(Baldassarre,, 2022, p. 252)\@BBCP. Due to their usefulness, intrinsically motivated RL methods have been studied extensively in the computational RL literature at large (see reviews by \@BBOPcitep\@BAP\@BBNColas et al.,, 2022; Aubret et al.,, 2023; Lidayan et al.,, 2024\@BBCP). An assumption underlying much work in RL is the reward hypothesis: “all of what we mean by goals and purposes can well be thought of as the maximization of the expected value of the cumulative sum of a received scalar signal (called reward)” (\@BBOPcitep\@BAP\@BBNSutton and Barto,, 2018, p. 53\@BBCP; cf., \@BBOPcitep\@BAP\@BBNSilver et al.,, 2021, p. 4\@BBCP). This hypothesis informs the formalisms used for competence-based intrinsic motivation (CB-IM), which tend to rely on the assumption that every goal can be sufficiently defined by a reward function. Intrinsically motivated RL is characterised by the use of task-agnostic intrinsic reward functions that collate or compare agent-internal variables independently of their semantics (\@BBOPcitep\@BAP\@BBNOudeyer and Kaplan,, 2008, p. 3\@BBCP; cf., \@BBOPcitep\@BAP\@BBNBerlyne,, 1965, p. 246\@BBCP and \@BBOPcitep\@BAP\@BBNOudeyer and Kaplan,, 2007\@BBCP).
@@
Goals can be viewed as “a set of constraints … that the agent seeks to respect” \@BBOPcitep\@BAP\@BBN(Colas et al.,, 2022, p. 1165, emphasis in original)\@BBCP. While the most immediate intuition of a goal is often as a desired state for the agent to reach (e.g., \@BBOPcitep\@BAP\@BBNKaelbling,, 1993, p. 1\@BBCP; \@BBOPcitep\@BAP\@BBNSchaul et al.,, 2015, p. 2\@BBCP), the formalism allows for a more general set of constraints on behaviour \@BBOPcitep\@BAP\@BBN(Lintunen et al., 2024b, , pp. 22–23)\@BBCP. In effect, any behaviour that can be defined by attempting to maximise some reward function on the environment can be formulated as a goal–skill pairing.
Aligning Computational Models With SDT
Based on a review of classic \@BBOPcitep\@BAP\@BBN(Oudeyer and Kaplan,, 2007)\@BBCP and state-of-the-art \@BBOPcitep\@BAP\@BBN(Colas et al.,, 2022; Aubret et al.,, 2023)\@BBCP computational IM, we next turn our focus to matching computational formalisms to each of the competence facets (Table 1). Our matching is non-exhaustive, and instead we focus on exemplifying how each facet could be formalised.
@@
Let us consider the case where competence in SDT refers to effectance, i.e., motivation by observing that one’s own actions cause change in the environment. To recognise effectance, an agent needs a mechanism to distinguish whether a change in the environment is caused by the agent itself or some external factor. An example of a computational model fitting effectance is Rewarding Impact-Driven Exploration (RIDE) \@BBOPcitep\@BAP\@BBN(Raileanu and Rocktäschel,, 2020)\@BBCP, as part of which the agent is rewarded for the amount of change it causes in its state observation: the “impact” of the agent’s actions, measured between two consecutive observations. Specifically, the impact-driven reward is defined as:
@@
In RIDE, the agent computes its reward using the learned representation of the observation as opposed to the raw sensory space (e.g., pixels). Inspired by \@BBOPcitet\@BAP\@BBNPathak et al., (2017)\@BBCP, the parameters of the model $\phi$ are optimised so that an agent learns the features it can control—that is, it learns to ignore aspects of the environment it cannot affect. Decomposing the reward function, the agent is rewarded for a combination of changes in those state features that it is able to consistently have some effect on (Eq. 1, numerator) and the novelty of the state into which its action leads (Eq. 1, denominator). The effect of the denominator is that the reward obtainable from particular states diminishes the more the agent observes them, even if the agent’s action has the same effective outcome as previously observed. In recent SDT texts, effectance motivation does not require novelty, but the original usage of the term required observations to be novel \@BBOPcitep\@BAP\@BBN(White,, 1959, p. 322)\@BBCP. By having the reward given by RIDE wane the more a particular situation is perceived, RIDE offers a simple formalisation of this original novelty aspect of effectance motivation.
@@
We next consider the case in which competence refers to skill use or the “exercise $\ldots$ of one’s capacities” \@BBOPcitep\@BAP\@BBN(Ryan and Deci,, 2017, p. 86)\@BBCP. Following \@BBOPcitet\@BAP\@BBNDeterding et al., (2024, p. 6)\@BBCP, we cover the specific understanding of “exercise” as implementing or bringing something to bear \@BBOPcitep\@BAP\@BBN(Merriam-Webster.com,, 2025)\@BBCP; an alternative interpretation as engaging a skill repeatedly in order to grow it will be treated under Capacity growth (C4). As per Table 1, competence as skill-use motivation can be understood in two different ways: observing an opportunity (an appropriate situation) to use a skill (C2.a) and that one uses a particular capacity in the course of one’s action (C2.b).
@@
Broadly, there are two sets of computational approaches used to model an agent learning about opportunities to use particular skills: those in which the agent ignores the current state of the world and those in which it does not. In both cases, hierarchy is used: an agent learns a distribution over its set of policies (i.e., a high-level policy) to determine which skill to use. In the former case, an agent follows the same policy regardless of its state observation—in this case, an “appropriate” opportunity can be thought of as a learning opportunity, determined not by observing the environment, but from information about how the learning is proceeding. In the latter case, identifying an opportunity to use a particular skill can also draw on observations of the environment.
@@
Specifically, the first case of C2.a, in which an agent is learning about opportunities to use a skill while ignoring the current state of the world, is well exemplified by computational approaches using Learning Progress (LP), which approximates some derivative of performance over time (initially proposed by \@BBOPcitep\@BAP\@BBNOudeyer et al.,, 2007, p. 7\@BBCP; see also related earlier work by \@BBOPcitep\@BAP\@BBNSchmidhuber,, 1991, p. 1461\@BBCP). In systems employing LP, it is typically used as a measure to support the selection of intermediate-difficulty goals with respect to the changing skills of an agent over time. LP, as an intrinsic reward for goal selection, has been well studied across both machine and human reinforcement learners (e.g., \@BBOPcitep\@BAP\@BBNColas et al.,, 2019; Ten et al.,, 2021; Molinaro et al.,, 2024\@BBCP), and could thus offer good starting points for formalising C2.a.
@@
While a reward function prioritises the choice of skills, an agent also needs a mechanism to select a skill from that prioritisation. One commonly used mechanism is Thompson sampling \@BBOPcitep\@BAP\@BBN(Thompson,, 1933)\@BBCP, also known as probability matching, where an agent draws a “belief”—choosing a skill according to a probability distribution (learned from some reward signal such as LP values) over its set of skills—and then acts greedily with respect to the drawn skill until it has more “evidence” (e.g., a change in LP values) and its “belief” (i.e., skill-selection policy) is updated. Thompson sampling has been shown to explain aspects of human exploration in certain tasks \@BBOPcitep\@BAP\@BBN(Gershman,, 2018)\@BBCP, making it a candidate for a psychologically plausible formalism of human skill selection.
@@
An example of a system formalising the second case of C2.a, in which an agent learns about opportunities to use particular skills while accounting for the current state of the world, is Variational Intrinsic Control (VIC). In VIC, the appropriate skill for a given starting state is decided using a separate learned model of how empowering the skill has been in the past. We mean empowering in the sense of its definition in the AI literature,111“Empowerment” has also been studied in the context of human exploration, e.g., as “exploring options that enable the generation of as many more options as possible” \@BBOPcitep\@BAP\@BBN(Brändle et al.,, 2023, p. 1482)\@BBCP. as the combined ability of an agent to control its environment and to sense this control (cf., \@BBOPcitep\@BAP\@BBNSalge et al.,, 2014, p. 2\@BBCP; \@BBOPcitep\@BAP\@BBNGregor et al.,, 2016, p. 2\@BBCP). The decision over skills is conditional on the state of the world, in that the policy for choosing skills is updated based on how empowering the skill turned out to be for that given situation. The distribution over skills is updated during skill use, such that more empowering skills become more likely to be chosen \@BBOPcitep\@BAP\@BBN(Gregor et al.,, 2016, p. 4)\@BBCP, using the reward \@BBOPcitep\@BAP\@BBN(Gregor et al.,, 2016, p. 2)\@BBCP:
$R_{t}({\mathbf{s}}_{0},{\mathbf{s}}_{t},{\mathbf{g}}):=\underbrace{\log q_{\bm{\phi}}({\mathbf{g}}\mid{\mathbf{s}}_{0},{\mathbf{s}}_{t})}_{\text{($\alpha$)}}\underbrace{-\log p_{\bm{\theta}}({\mathbf{g}}\mid{\mathbf{s}}_{0})}_{\text{($\beta$)}}.$ (2)
@@
The discriminator, $q_{\bm{\phi}}$, is an arbitrary variational distribution parametrised by ${\bm{\phi}}$@BBOPcitep\@BAP\@BBN(Barber and Agakov,, 2003, p. 2)\@BBCP. Given the first (${\mathbf{s}}_{0}$) and last (${\mathbf{s}}_{t}$) observations induced by the skill corresponding to the goal being pursued, $q_{\bm{\phi}}$ defines a probability distribution over goals. Successfully discriminating goals in the observation space requires the agent to observe distinct regions of the state space. If a goal is not discriminable based on observations, two or more skills are producing overlapping behaviours (and therefore, the skills lack diversity); conversely, if a goal is discriminable, then the corresponding skill is inducing trajectories unique to that skill. Thus, this reward term is maximised when there is no uncertainty in the prediction.
@@
Given a starting state, ${\mathbf{s}}_{0}$, a goal, ${\mathbf{g}}$, is selected according to the distribution defined by $q_{\bm{\phi}}$, which the agent pursues until the termination state ${\mathbf{s}}_{t}$. Then, $p_{\bm{\theta}}$ is reinforced based on how empowering the corresponding skill was (estimated using Eq. 2). The final state becomes the new starting state, ${\mathbf{s}}_{0}\leftarrow{\mathbf{s}}_{f}$, and the process is repeated. This can motivate the agent to follow skills that lead to distinct end states ($\alpha$) and learn a wide range of skills for any given situation ($\beta$).
@@
While C2.a is more about finding suitable opportunities to use particular skills, such as in the case of VIC learning a state-conditional policy over skills, C2.b is more about an agent being motivated by observing that a skill is in use at all. The first term of the VIC reward function (Eq. 2, $\alpha$) can be adapted to emphasise an agent’s ability to distinguish that a particular skill is in use at a given time, by only conditioning on the current state. The reward for one such system, Diversity Is All You Need (DIAYN) \@BBOPcitep\@BAP\@BBN(Eysenbach et al.,, 2019)\@BBCP, can be defined [^2] See \@BBOPcitep\@BAP\@BBNEysenbach et al.,, 2019, Appendix A, p. 12\@BBCP for discussion on a constant term omitted here. as:
@@
where $q_{\bm{\phi}}$ is an arbitrary variational distribution parametrised by ${\bm{\phi}}$ (refer back to the text following Eq. 2 for detail). An agent is rewarded for its ability to distinguish the current goal, ${\mathbf{g}}$; thus, the overall return is maximised when the agent can perfectly infer which of its skills is in use at any given time.
DIAYN, in further differentiation from VIC, ignores the state of the world in its choice of skill, instead choosing skills uniformly at random—in other words, learning skills that are empowering on average over the state space. For further examples, one can turn to other variational empowerment methods (see \@BBOPcitep\@BAP\@BBNChoi et al.,, 2021, p. 3\@BBCP). The choice to fix the distribution over skills as uniform—although not an intuitive choice for a psychologically plausible model of goal selection, except maybe in special cases such as the absence of prior information—is common in the CB-IM literature. This approach ensures that all skills receive, in expectation, an equal amount of training signal to improve. DIAYN learns a diverse set of skills across a wide array of situations, such as running, walking, hopping, flipping, or gliding to move through a space \@BBOPcitep\@BAP\@BBN(Eysenbach et al.,, 2019, p. 5)\@BBCP. It thus shows that a formalism motivated by skill use can drive skill learning.

### v2

@@ Abstract
Computational modelling offers a powerful tool for formalising psychological theories, making them more transparent, testable, and applicable in digital contexts. Yet, the question often remains: how should one computationally model a theory? We provide a demonstration of how formalisms taken from artificial intelligence can offer a fertile starting point. Specifically, we focus on the “need for competence”, postulated as a key basic psychological need within Self-Determination Theory (SDT)—arguably the most influential framework for intrinsic motivation (IM) in psychology. Recent research has identified multiple distinct facets of competence in key SDT texts: effectance, skill use, task performance, and capacity growth. We draw on the computational IM literature in reinforcement learning to suggest that different existing formalisms may be appropriate for modelling these different facets. Using these formalisms, we reveal underlying preconditions that SDT fails to make explicit, demonstrating how computational models can improve our understanding of IM. More generally, our work can support a cycle of theory development by inspiring new computational models, which can then be tested empirically to refine the theory. Thus, we provide a foundation for advancing competence-related theory in SDT and motivational psychology more broadly.
Keywords: need for competence; self-determination theory; intrinsic motivation; theory development; intrinsically motivated reinforcement learning; computational modelling
@@ Introduction
Self-Determination Theory (SDT) \@BBOPcitep\@BAP\@BBN(Deci and Ryan,, 1985)\@BBCP is one of the most influential psychological theories of motivation, specifically intrinsic motivation (IM). Notably, SDT, like many psychological theories, is expressed in “soft” verbal propositions. In light of the current theory crisis \@BBOPcitep\@BAP\@BBN(Oberauer and Lewandowsky,, 2019)\@BBCP, only-verbal theory has drawn increasing critique: it can leave many auxiliary assumptions and boundary conditions unspecified, making it hard to derive hypotheses that can be rigorously tested \@BBOPcitep\@BAP\@BBN(Scheel et al.,, 2021)\@BBCP. Many psychology researchers have advocated for computational modelling as a paradigm (e.g., \@BBOPcitep\@BAP\@BBNMarsella et al.,, 2010; Oberauer and Lewandowsky,, 2019; Robinaugh et al.,, 2021\@BBCP) or even as a necessity \@BBOPcitep\@BAP\@BBN(Van Dongen et al.,, 2024)\@BBCP for theory development. Yet formal specification as computational modelling remains largely absent in SDT, judging by, e.g., the most recent integrative volume of the theory \@BBOPcitep\@BAP\@BBN(Ryan,, 2023)\@BBCP.
A major challenge of computational modelling is inferring plausible formalisms for a given theory. We make the case that computational IM, i.e., formalisms developed to drive the behaviour of artificial agents, can inform the development of SDT theory in human motivation research. Computational IM has been informed by SDT (e.g., \@BBOPcitep\@BAP\@BBNOudeyer and Kaplan,, 2007\@BBCP), but has undergone major advancements independent from human psychology, addressing problems such as open-ended development, task-agnostic learning, and dealing with the potential sparsity of rewards \@BBOPcitep\@BAP\@BBN(Colas et al.,, 2022, p. 1161)\@BBCP.
Our case study focuses on the need for competence, one of the basic psychological needs underpinning IM in SDT. Through conceptual analysis \@BBOPcitep\@BAP\@BBN(Machado and Silva,, 2007)\@BBCP, \@BBOPcitet\@BAP\@BBNDeterding et al., (2024)\@BBCP identify four distinct facets of competence in canonic SDT texts, reproduced below in Table 1.
@@
Task performance (C3) Observing that one performs (a) well at an intended task and/or (b) to an extent that requires a certain skill or skill level.
@@
In the section Matching Computational Models With SDT, we match each distinct facet with examples of formalisms from the computational IM literature and discuss how each example formalises aspects of the theory. These formalisms not only lay a foundation for computationally specifying the need for competence, but, by linking to the corresponding papers, we also unveil a wealth of existing implementations, simulators, and evaluation metrics that could be used in psychological motivation research. The value of such resources is detailed in the section Future Work and Conclusion, but, for one, they empower future testing of whether formalisms fitting the facets produce empirical data aligned with propositions from the theory.
@@
We begin by justifying why reinforcement learning (RL) is a suitable domain from which to draw computational models of competence. In addition, to support understanding of the models we discuss, we introduce key concepts from intrinsically motivated RL—notably, goals and skills and their relationship to the reward function—and point to further reading.
The idea that IM might be driven by biological reward systems (\@BBOPcitep\@BAP\@BBNBarto et al.,, 2004, p. 113\@BBCP, in reference to \@BBOPcitep\@BAP\@BBNKakade and Dayan,, 2002; Dayan and Balleine,, 2002\@BBCP) has driven the choice of methods centrally used in computational IM: RL \@BBOPcitep\@BAP\@BBN(Baldassarre,, 2022, p. 252)\@BBCP. Due to their usefulness, intrinsically motivated RL methods have been studied extensively in the computational RL literature at large (see reviews by \@BBOPcitep\@BAP\@BBNColas et al.,, 2022; Aubret et al.,, 2023; Lidayan et al.,, 2024\@BBCP). Central to RL is the construct of reward function, and an assumption underlying much work in RL is the reward hypothesis: “all of what we mean by goals and purposes can well be thought of as the maximization of the expected value of the cumulative sum of a received scalar signal (called reward)” (\@BBOPcitep\@BAP\@BBNSutton and Barto,, 2018, p. 53\@BBCP; cf., \@BBOPcitep\@BAP\@BBNSilver et al.,, 2021, p. 4\@BBCP). This hypothesis informs the formalisms used for competence-based intrinsic motivation (CB-IM), which tend to rely on the assumption that every goal can be sufficiently defined by a reward function. Intrinsically motivated RL is characterised by the use of task-agnostic intrinsic reward functions that collate or compare agent-internal variables independently of their semantics (\@BBOPcitep\@BAP\@BBNOudeyer and Kaplan,, 2008, p. 3\@BBCP; cf., \@BBOPcitep\@BAP\@BBNBerlyne,, 1965, p. 246\@BBCP and \@BBOPcitep\@BAP\@BBNOudeyer and Kaplan,, 2007\@BBCP).
@@
Goals can be viewed as “a set of constraints … that the agent seeks to respect” \@BBOPcitep\@BAP\@BBN(Colas et al.,, 2022, p. 1165, emphasis in original)\@BBCP. While the most immediate intuition of a goal is often as a desired state for the agent to reach (e.g., \@BBOPcitep\@BAP\@BBNKaelbling,, 1993, p. 1\@BBCP; \@BBOPcitep\@BAP\@BBNSchaul et al.,, 2015, p. 2\@BBCP), the formalism allows for a more general set of constraints on behaviour. In effect, any behaviour that can be defined by attempting to maximise some reward function on the environment can be formulated as a goal–skill pairing.
Matching Computational Models With SDT
We next turn our focus to matching computational formalisms to each of the competence facets in Table 1 by providing examples from RL. The formalisms have been identified through review of both classic \@BBOPcitep\@BAP\@BBN(Oudeyer and Kaplan,, 2007)\@BBCP and state-of-the-art \@BBOPcitep\@BAP\@BBN(Colas et al.,, 2022; Aubret et al.,, 2023)\@BBCP computational IM. An exhaustive matching is beyond the scope of this paper. Rather, we highlight formalisms that we believe to fit the verbal definitions of each facet and reveal preconditions that SDT implicitly assumes. We focus on intuition about how each model fits the theory. For more intuition and detail on how each model is designed, implemented, and evaluated, we direct the reader to the corresponding papers.
@@
Let us consider the case where competence in SDT refers to effectance, i.e., motivation by observing that one’s own actions cause change in the environment. To recognise effectance, an agent needs a mechanism to distinguish whether a change in the environment is caused by the agent itself or some external factor. An example of a computational model fitting effectance is Rewarding Impact-Driven Exploration (RIDE) \@BBOPcitep\@BAP\@BBN(Raileanu and Rocktäschel,, 2020)\@BBCP, as part of which the agent is rewarded for the amount of change it causes via a single action: the “impact” of the agent’s action, measured between two consecutive observations. Specifically, the impact-driven reward is defined as:
@@
In RIDE, the agent computes its reward using the learned representation of the observation ($\phi({\mathbf{s}}_{t})$) as opposed to the raw sensory space (e.g., pixels). The parameters of the representation model $\phi$ are optimised so that an agent only learns the features it can control—that is, it learns to ignore aspects of the environment it cannot affect (inspired by \@BBOPcitep\@BAP\@BBNPathak et al.,, 2017\@BBCP). The agent is rewarded for a combination of changes in only those state features that its own actions are able to consistently have some effect on (Eq. 1, numerator). This reward is scaled according to the novelty of the state into which its action leads (Eq. 1, denominator). The effect of the denominator is that the reward obtainable from particular states diminishes the more the agent observes them, even if the agent’s action has the same effective outcome as previously observed. In recent SDT texts, effectance motivation does not require novelty, but the original usage of the term required observations to be novel \@BBOPcitep\@BAP\@BBN(White,, 1959, p. 322)\@BBCP. By having the reward wane the more a particular situation is perceived, RIDE offers a simple formalisation of this original novelty aspect of effectance motivation, while ensuring the agent is only rewarded for changes its own actions have caused.
@@
We next consider the case in which competence refers to skill use or the “exercise $\ldots$ of one’s capacities” \@BBOPcitep\@BAP\@BBN(Ryan and Deci,, 2017, p. 86)\@BBCP. Following \@BBOPcitet\@BAP\@BBNDeterding et al., (2024, p. 6)\@BBCP, we cover the specific understanding of “exercise” as implementing or bringing something to bear \@BBOPcitep\@BAP\@BBN(Merriam-Webster.com,, 2025)\@BBCP. [^1] An alternative interpretation of “exercise”, as engaging a skill repeatedly to grow it, will be treated under Capacity growth (C4). As per Table 1, competence as skill-use motivation can be understood in two different ways: observing an opportunity (an appropriate situation) to use a skill (C2.a) or that one uses a particular capacity in the course of one’s action (C2.b).
@@
Broadly, there are two kinds of computational approaches used to model an agent learning about opportunities to use particular skills: those in which the agent ignores the current state of the world and those in which it does not. In both cases the agent learns a policy over its set of skills to determine which skill to use. In the former case, an agent follows the same policy regardless of its state observation—in this case, an “appropriate” opportunity can be thought of as a learning opportunity, determined not based on the observation of the environment, but on how the agent’s learning is proceeding. In the latter case, identifying an opportunity to use a particular skill can also depend on the observation.
@@
Specifically, the first case of C2.a, in which an agent is learning about opportunities to use a skill while ignoring the current state of the world, is well exemplified by computational approaches using Learning Progress (LP), which approximates some derivative of performance over time (initially proposed by \@BBOPcitep\@BAP\@BBNOudeyer et al.,, 2007, p. 7\@BBCP; see also related earlier work by \@BBOPcitep\@BAP\@BBNSchmidhuber,, 1991, p. 1461\@BBCP). In systems employing LP, it is typically used as a measure to support selecting intermediate-difficulty goals with respect to the changing skills of an agent over time. In fact, LP, as an intrinsic reward for goal selection, has been well studied across both machine and human reinforcement learners (e.g., \@BBOPcitep\@BAP\@BBNColas et al.,, 2019; Ten et al.,, 2021; Molinaro et al.,, 2024\@BBCP), and could thus offer good starting points for formalising C2.a.
@@
While a (high-level) reward function prioritises the choice of skills, an agent also needs a mechanism to select a skill from that prioritisation. One commonly used mechanism is Thompson sampling \@BBOPcitep\@BAP\@BBN(Thompson,, 1933)\@BBCP, also known as probability matching, in which an agent draws a skill according to a “belief”—a probability distribution (learned from a reward signal such as LP values) over its set of skills—and then acts greedily with respect to the drawn skill until it has more “evidence” (e.g., a change in LP values) and its “belief” (i.e., skill-selection policy) is updated. Thompson sampling has been shown to explain aspects of human exploration in certain tasks \@BBOPcitep\@BAP\@BBN(Gershman,, 2018)\@BBCP, making it a candidate for a psychologically plausible formalism of human skill selection.
@@
An example of a system formalising the second case of C2.a, in which an agent learns about opportunities to use particular skills while accounting for the current state of the world, is Variational Intrinsic Control (VIC). In VIC, the appropriate skill for a given starting state is decided using a separate learned distribution dependant on how empowering the skill has been in the past. We mean empowering in the sense of its definition in the artificial intelligence (AI) literature,222“Empowerment” has also been studied in the context of human exploration, e.g., as “exploring options that enable the generation of as many more options as possible” \@BBOPcitep\@BAP\@BBN(Brändle et al.,, 2023, p. 1482)\@BBCP. as the combined ability of an agent to control its environment and to sense this control (cf., \@BBOPcitep\@BAP\@BBNSalge et al.,, 2014, p. 2\@BBCP; \@BBOPcitep\@BAP\@BBNGregor et al.,, 2016, p. 2\@BBCP). The decision over skills is conditional on the state of the world, in that the policy for choosing skills is updated based on how empowering the skill turned out to be for that given situation. The distribution over skills is updated after skill use, such that more empowering skills become more likely to be chosen \@BBOPcitep\@BAP\@BBN(Gregor et al.,, 2016, p. 4)\@BBCP, using the reward \@BBOPcitep\@BAP\@BBN(Gregor et al.,, 2016, p. 2)\@BBCP:
$R_{t}({\mathbf{s}}_{0},{\mathbf{s}}_{f},{\mathbf{g}}):=\underbrace{\log q_{\bm{\phi}}({\mathbf{g}}\mid{\mathbf{s}}_{0},{\mathbf{s}}_{f})}_{\text{($\alpha$)}}\underbrace{-\log p_{\bm{\theta}}({\mathbf{g}}\mid{\mathbf{s}}_{0})}_{\text{($\beta$)}}.$ (2)
@@
The discriminator, $q_{\bm{\phi}}$, is an arbitrary variational distribution parametrised by ${\bm{\phi}}$@BBOPcitep\@BAP\@BBN(Barber and Agakov,, 2003, p. 2)\@BBCP. Given the first (${\mathbf{s}}_{0}$) and last (${\mathbf{s}}_{f}$) observations induced by the skill corresponding to the goal, ${\mathbf{g}}$, being pursued, $q_{\bm{\phi}}$ defines a probability distribution over goals. Discriminating goals in the observation space requires the agent to observe distinct regions of the observation space. If a goal is not discriminable based on observations, two or more skills are producing overlapping behaviours; conversely, if a goal is discriminable, then the corresponding skill is inducing trajectories unique to that skill. Thus, this reward term is maximised when there is no uncertainty in the prediction.
@@
Given a starting state, ${\mathbf{s}}_{0}$, a goal, ${\mathbf{g}}$, is selected according to the distribution defined by $p_{\bm{\theta}}$, which the agent pursues until the final state ${\mathbf{s}}_{f}$. Then, $p_{\bm{\theta}}$ is reinforced based on how empowering the corresponding skill was (estimated using Eq. 2). The final state becomes the new starting state, ${\mathbf{s}}_{0}\leftarrow{\mathbf{s}}_{f}$, and the process is repeated. This can motivate the agent to ($\alpha$) follow skills that lead to distinct final states and ($\beta$) learn a wide range of skills for any given situation. Learning a state-conditional policy over skills fits the understanding of competence as finding suitable opportunities to use particular skills.
@@
While C2.a is about finding suitable opportunities to use particular skills, C2.b is more about an agent being motivated by observing that a skill is in use at all. The first term of the VIC reward function (Eq. 2, $\alpha$) can be adapted to emphasise an agent’s ability to distinguish that a particular skill is in use, by only conditioning on the current state. The reward for one such system, Diversity Is All You Need (DIAYN) \@BBOPcitep\@BAP\@BBN(Eysenbach et al.,, 2019)\@BBCP, can be defined [^3] See \@BBOPcitep\@BAP\@BBNEysenbach et al.,, 2019, Appendix A, p. 12\@BBCP for discussion on a constant term omitted here. as:
@@
where $q_{\bm{\phi}}$ is a discriminator parametrised by ${\bm{\phi}}$ (refer back to the text following Eq. 2 for detail). An agent is rewarded for its ability to distinguish the current goal, ${\mathbf{g}}$; thus, the overall return is maximised when the agent can perfectly infer which of its skills is in use at any given time.
DIAYN, in further differentiation from VIC, ignores the state of the world in its choice of skill—instead, choosing skills uniformly at random, thus learning skills that are empowering on average over the observation space. Fixing the distribution over skills as uniform, although not an intuitive choice for a psychologically plausible model of goal selection (except maybe in cases such as the absence of prior information) is common in the CB-IM literature. This approach ensures that all skills receive, in expectation, an equal amount of training signal to improve. DIAYN learns a diverse set of skills—running, walking, hopping, flipping, and gliding—across a wide array of situations \@BBOPcitep\@BAP\@BBN(Eysenbach et al.,, 2019, p. 5)\@BBCP. It thus shows that recognising which skill is in use, that is, being able to distinguish the use of a particular capacity in the course of one’s action, can drive skill learning.
Beyond DIAYN and VIC, one can turn to other variational empowerment methods \@BBOPcitep\@BAP\@BBN(see Choi et al.,, 2021)\@BBCP for further examples of CB-IMs fitting competence as skill use.